OpenStudy (anonymous):

Determine algebraically whether the function is even, odd or neither. f(x)=x^1/3

OpenStudy (anonymous):

definitely odd

OpenStudy (anonymous):

since \[f(-3)=\sqrt[3]{-x}=-\sqrt[3]{x}=-f(x)\] for example \[f(-8)=\sqrt[3]{-8}=-2=-\sqrt[3]{8}\]